OpenStudy (anonymous):

Help Me Please! Divide. a^2+5a+4/a^3 divide a^2+3a+2/a^2-2a

OpenStudy (anonymous):

so far I have \[a+4/a^3\div a+2/a(a-2)\] but I have no clue if that's right or even close...

OpenStudy (anonymous):

Well, the first thing to remember is that dividing is the same as multiplying by the reciprocal (flip). So, you're trying to simplify: \[\frac{a^2+5a+4}{a^3} \times \frac{a^2−2a}{a^2+3a+2}\] See how the term on the right is what you said but flipped over? OK, now you need to factor all the terms you can, to see what will cancel: \[\frac{(a+1)(a+4)}{a^3} \times \frac{a(a−2)}{(a+2)(a+1)}\] Now, can you see terms you can cancel? Remember, with fractions, if the same thing is on the top and bottom of a fraction, it cancels out.

OpenStudy (anonymous):

And if you look at yours... looks like you're right! (just not as nicely formatted). I'd change that divide into a multiply - divides are always confusing and completely unnecessary since we know what they _really_ are.

OpenStudy (anonymous):

thank you so much! @gbell

OpenStudy (anonymous):

can you help me with one more? @gbell

OpenStudy (anonymous):

Sure! Post a new question (not on this one) then send me a message.